OpenStudy (anonymous):

Sue is in an airplane which travels at a constant 300 kilometers per hour. The angle of climb of the airplane is 30 ∘ . How long (in seconds) would it take from takeoff, before the airplane reaches 3000 meters, when Sue can access Wifi and work on the rest of the weekly problems?

OpenStudy (anonymous):

The distance traveled is equal to the velocity times the time travelled. So, lets convert the airplane's speed to m/s: \[300 km/h \times \frac{ .277777 m/s }{ 1 km/h } = 83.33333 m/s\]So, after t seconds, the plane will travel 83.33333t meters. Now lets make a picture using this distance:  Do you know how to solve from here?

OpenStudy (anonymous):

Unfortunately i do not :( I just wanted to know the answer.

OpenStudy (anonymous):

DO you use a trig function

OpenStudy (anonymous):

Yes, you could use sine, since we have the side opposite the angle and the hypotenuse, or we can use the law of sines

OpenStudy (anonymous):

sin\[\sin30=3000/83.3333\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

Yes, but you left out the variable, which should go with the 83

OpenStudy (anonymous):

\[\sin 30 = 3000 / (83.33333t)\]

OpenStudy (anonymous):

\[t=3000/\sin30(83.333)\]

OpenStudy (anonymous):

Exactly, so what are you getting as an answer?

OpenStudy (anonymous):

i dont' have a calculator with me

OpenStudy (anonymous):

Ok, I'm getting 72 seconds

OpenStudy (anonymous):

that's correct